OpenStudy (anonymous):

Solve the inequality x^2 + 4x >= 32

OpenStudy (anonymous):

x2+4x-32>=0 x2+8x-4x-32>=0 x(x+8)-4(x+8)>=0 (x+8)(x-4)>=0 [x-(-8)] (x-4)>=0 implies x<= -8 and x>=4